Can humans choose a number randomly? Fifty adults were asked to chose a random number from one to eight. Use α = 0.05 as the level of significance. The table below summarizes the results.
Number 1 2 3 4 5 6 7 8 Frequency 5 3 10 13 7 2 4 6

(a) Calculate the degrees of freedom

(b)Calculate the test statistic x^2

(c)If the critical value is χ2 0.05 =14.067 should you accept or reject the null hypothesis?
Answer #1

(a) Calculate the degrees of freedom

As n = 8

the df = n - 1 = 8 - 1 = 7

df = 7

(b)Calculate the test statistic x^2

Event Observed Frequency(O) Expected Probability Expected Frequency ( E) ( O-E)^2 ( O-E)^2/E
1 5 0.125 6.25 1.5625 0.25
2 3 0.125 6.25 10.5625 1.69
3 10 0.125 6.25 14.0625 2.25
4 13 0.125 6.25 45.5625 7.29
5 7 0.125 6.25 0.5625 0.09
6 2 0.125 6.25 18.0625 2.89
7 4 0.125 6.25 5.0625 0.81
8 6 0.125 6.25 0.0625 0.01
Sum 50 1 50 - 15.28

\chi^2 = 15.28

(c)If the critical value is χ2 0.05 =14.067 should you accept or reject the null hypothesis?

Null hypothesis: The probability are equal

Alternate hypothesis: At least one probability is different

We reject the null hypothesis

If \chi^2 > χ2 0.05 ( 15.28 > 14.067 )

hence the null hypothesis is rejected